### v4

# Online Learning of Nonlinear Parametric Models under Non-smooth Regularization using EKF and ADMM

Abstract
This paper proposes a novel combination of extended Kalman filtering (EKF) with the alternating direction method of multipliers (ADMM) for learning parametric nonlinear models online under non-smooth regularization terms, including $\ell_{1}$ and $\ell_{0}$ penalties and bound constraints on model parameters. For the case of linear time-varying models and non-smooth convex regularization terms, we provide a sublinear regret bound that ensures the proper behavior of the online learning strategy. The approach is computationally efficient for a wide range of regularization terms, which makes it appealing for its use in embedded control applications for online model adaptation. We show the performance of the proposed method in three simulation examples, highlighting its effectiveness compared to other batch and online algorithms.
keywords:
Kalman filtering; non-smooth regularization; online learning; parameter estimation; adaptive control; neural networks.
thanks: The research work of Lapo Frascati has been financially supported by ODYS S.r.l. The research work of Alberto Bemporad has been funded by the European Union (ERC Advanced Research Grant COMPACT, No. 101141351). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
1 Introduction
Online learning of nonlinear parametric models is of paramount importance in several domains, including model-based adaptive control and real-time estimation of unmeasured variables. Typically, parametric models derived from physics [bib9] or black-box [bib8] structures are identified offline on training data, then directly deployed and used without any further updates. On the other hand, further adapting the model online can significantly improve its predictive capabilities [bib10], especially when the phenomenon we are modeling changes over time, and allows for smaller model structures that adapt to varying operating conditions, unlike single, overall models trained offline to cover all conditions.
A vast literature exists for online learning [bib7], and several approaches using first or second order gradient information are suitable for real-time model adaptation. Examples of algorithms exploiting first-order information are online gradient descent (OGD) and its regularized alternatives, i.e., follow the regularized leader (FTRL) and online mirror descent (OMD) [bib20], which are usually fast at execution but slow at convergence. Second-order gradient information can be used to improve the convergence speed, such as in the online Newton step (ONS), or in the well known extended Kalman filter (EKF) [bib14] which has been proved to be very effective in online model adaptation by treating the parameters as constant states to be estimated [bib1, bib3].
While FTRL and OMD, as well as modified versions of the standard EKF [bib1], can effectively deal with smooth regularization terms, non-smooth regularizers are often required when learning models. Non-smooth $\ell_{0}$ or $\ell_{1}$ penalties are used to induce sparsity in the model [bib25], and group-Lasso terms to remove entire parts of the model, such as neurons of a neural network [bib26]. Indeed, obtaining compact models is particularly important for embedded applications, such as model predictive control, where the real-time numerical complexity of the controller depends directly on the complexity of the prediction model. In addition, indicator functions of feasible sets are another example of non-smooth penalties that are used to impose constraints on model coefficients, such as known bounds on certain unknown physical parameters [bib5].
Several approaches have been proposed for online learning under non-smooth regularization. Online ADMM (alternating direction method of multipliers) [bib2] allows handling quite general non-smooth regularizers, while EKF can be modified to deal with $\ell_{1}$-regularization by treating it as a special limit case of a smooth regularizer [bib1]. While the main limitation of online ADMM is its convergence speed, the main limitation of EKF is that it is not directly suitable for dealing with general non-smooth regularization terms, such as $\ell_{0}$ regularization, group-Lasso penalties, and indicator functions of feasible sets. The main contribution of this work is to show interesting connections and similarities between these two approaches, developing an extension of EKF that inherits the advantages from both, i.e., the convergence speed of EKF and the capability to handle general non-smooth regularization terms of online-ADMM.
The proposed method consists of a simple and computationally efficient modification of the EKF algorithm by intertwining updates based on online measurements and output prediction errors with updates related to ADMM iterations. This modification allows EKF to deal with a broad class of non-smooth regularization terms for which ADMM is applicable, including $\ell_{0}$$\ell_{1}$ penalties, group-Lasso and indicator functions of simple sets. For linear time-varying models and convex regularization terms, we provide a sublinear regret bound that proves the proper behavior of the resulting online learning strategy. The proposed method is computationally efficient and numerically robust, making it especially appealing for embedded adaptive control applications.
The rest of the paper is organized as follows. Section 2 gives a quick introduction to the use of EKF for online model learning, setting the background for the proposed ADMM+EKF approach described in Section 3. In Section 4, we prove a sublinear regret bound for the proposed approach in the convex linear case. Simulation results are shown in Section 5 and conclusions are drawn in Section 6.
1.1 Notation
Given a matrix $A\in\mathbb{R}^{m\times n}$, $A_{i:}$ denotes the ith row of A, $A_{:j}$ its $j$th column, $A_{ij}$ its $(i,j)$th entry. Given a vector $v\in\mathbb{R}^{n}$ we denote by $v^{m}$ a measurement of $v$, by $\lVert v\rVert_{1}=\sum_{i=1}^{n}|v_{i}|$ the 1-norm of $v$ and by $\lVert v\rVert_{0}$ its 0-norm, which is defined as the number of non-zero elements in the vector. Given a symmetric positive semidefinite matrix $P=P^{T}\succeq 0$, $P\in\mathbb{R}^{n\times n}$, we denote by $\lVert v\rVert^{2}_{P}$ the quadratic form $v^{T}Pv$. Further, we denote by $\hat{v}_{i|j}$ the estimate of vector $v$ at instant $i$ given all information up to instant $j$, and by $P_{i|j}$ the corresponding covariance matrix. $v\sim\mathcal{N}(\mu,P)$ and $v\sim\mathcal{U}_{[v_{min},v_{max}]}$ denote that v was randomly generated from a normal distribution with mean $\mu$ and covariance $P$ or from a uniform distribution in the interval $v_{min},v_{max}]$, respectively.
2 EKF for online model learning
Given a dataset $(z_{k}^{m},y_{k}^{m})$, $z\in\mathbb{R}^{n_{z}}$, $y\in\mathbb{R}^{n_{y}}$, $k=0,1,\ldots,N-1$, our goal is to recursively estimate a nonlinear parametric model
$y=h(k,z;x)$ (1)
which describes the (possibly time-varying) relationship between the input signal $z_{k}^{m}$ and the output signal $y_{k}^{m}$. In (1), $x\in\mathbb{R}^{n_{x}}$ is the vector of parameters to be learned, such as the weights of a feedforward neural network mapping $z$ into $y$, or the coefficients of a nonlinear autoregressive model, with $y$ representing the current output and $z$ a vector of past inputs and outputs of a dynamical system. In order to estimate $x$ and capture its possible time-varying nature, we consider the nonlinear dynamical model
$x_{k+1}=x_{k}+q_{k},\quad y_{k}^{nl}=h_{k}(x_{k})+r_{k}$ (2)
where $h_{k}(\cdot)=h(k,z_{k};\cdot)$ and we assume $h_{k}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{y}}$ differentiable for all $k$, $y_{k}^{nl}\in\mathbb{R}^{n_{y}}$, and $r_{k}\sim\mathcal{N}(0,R_{k}),q_{k}\sim\mathcal{N}(0,Q_{k})$ are the measurement and process noise that we introduce to model, respectively, measurement errors and variations of the model parameters over time, with corresponding covariance matrices $R_{k}=R_{k}^{\prime}\succ 0$, $Q_{k}=Q_{k}^{\prime}\succ 0$. By linearizing model (2) around a value $\overline{x}_{k}$ of the parameter vector, i.e., by approximating $h_{k}(x_{k})\approx h_{k}(\overline{x}_{k})+C_{k}(x_{k}-\overline{x}_{k})$, $C_{k,i:}=\nabla_{x}h_{ki}(\overline{x}_{k})^{\prime}$, $i=1,\ldots,n_{y}$, we obtain the linear time-varying model
$x_{k+1}=x_{k}+q_{k},\quad y_{k}=C_{k}x_{k}+r_{k}$ (3)
with $y_{k}=y_{k}^{nl}-h_{k}(\overline{x}_{k})+C_{k}\overline{x}_{k}$. The classical Kalman filter [bib14] can be used to estimate the state in (3), i.e., to learn the parameters $x_{k}$ recursively. Given $\hat{x}_{0|-1},P_{0|-1}$ we perform the following iterations for $k=0,\dots,N-1$:
$P_{k|k}^{-1}=P_{k|k-1}^{-1}+C_{k}^{T}R_{k}^{-1}C_{k}$ (4)
$\hat{x}_{k|k}=\hat{x}_{k|k-1}+P_{k|k}C_{k}^{T}R_{k}^{-1}(y_{k}^{m}-C_{k}\hat{x}_{k|k-1})$
$P_{k+1|k}^{-1}=(Q_{k}+P_{k|k})^{-1}$
$\hat{x}_{k+1|k}=\hat{x}_{k|k}$
with $\overline{x}_{k}=\hat{x}_{k|k-1}$. The first two updates in (4) are usually referred to as the correction step and the last two as the prediction step. Note that (4) is an EKF for model (2), since $C_{k}$ is the Jacobian of the output function at $\hat{x}_{k|k-1}$ and the output prediction error used in the correction step is $e_{k}=y_{k}^{m}-C_{k}\hat{x}_{k|k-1}=y_{k}^{nl,m}-h_{k}(\hat{x}_{k|k-1})$.
As shown in [bib13], the state estimates $\hat{x}_{k|k},\hat{x}_{k+1|k}$ generated by the Kalman filter (4) are part of the optimizer of the following optimization problem
$\hat{x}_{0|k},\dots,\hat{x}_{k|k},\hat{x}_{k+1|k}=\operatorname*{arg\,min}_{x_{0},\dots,x_{k},x_{k+1}}\left\lVert x_{0}-\hat{x}_{0|-1}\right\rVert^{2}_{P_{0|-1}^{-1}}+$ (5)
$+\sum_{i=0}^{k}\left\lVert y_{i}^{m}-C_{i}x_{i}\right\rVert^{2}_{R_{i}^{-1}}+\left\lVert x_{i+1}-x_{i}\right\rVert^{2}_{Q_{i}^{-1}}.$
Problem (5) can be solved recursively at each step $k$ by minimizing the following cost functions:
$\hat{x}_{k|k}=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}\hskip-10.00002pt+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}$
$\hat{x}_{k|k},\hat{x}_{k+1|k}=\operatorname*{arg\,min}_{x_{k},x_{k+1}}\left\lVert x_{k}-\hat{x}_{k|k}\right\rVert^{2}_{P_{k|k}^{-1}}+$
$\hskip 120.00018pt+\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{Q_{k}^{-1}}$
where $\hat{x}_{k|k},\hat{x}_{k+1|k},P_{k|k}$, and $P_{k+1|k}$ are the state estimates and covariance matrices computed as in (4).
3 EKF under non-smooth regularization
In order to regularize the model, we modify the classical iterations (4) by changing the minimization in (6b) to
$\hat{x}_{k|k}=$ $\operatorname*{arg\,min}_{x_{k}}\frac{1}{2}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+$ (7)
$+\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+g(x_{k})$
where $g(\cdot):\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}\cup\{+\infty\}$ is a possibly non-smooth and non-convex regularization term. By defining $\mathcal{S}=\{(x_{k},\nu)\in\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{x}}:\ x_{k}=\nu\}$, (7) can be equivalently reformulated as the following constrained optimization problem
$\hat{x}_{k|k},\nu^{\star}=$ $\operatorname*{arg\,min}_{(x_{k},\nu)\in\mathcal{S}}\frac{1}{2}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+$ (8)
$+\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+g(\nu)$
which can be solved by executing the following scaled ADMM iterations [bib5]:
$\hat{x}_{k|k}^{t+1}$ $=$ $\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+$ (9a)
$+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+\rho\left\lVert x_{k}-\nu^{t}+w^{t}\right\rVert_{2}^{2}$
$\nu^{t+1}$ $=$ $\operatorname*{arg\,min}_{\nu}g(\nu)+\frac{\rho}{2}\left\lVert\nu-\hat{x}_{k|k}^{t+1}-w^{l}\right\rVert_{2}^{2}$ (9b)
$=$ $\text{prox}_{\frac{g}{\rho}}(\hat{x}_{k|k}^{t+1}+w^{t})$
$w^{t+1}$ $=$ $w^{t}+\hat{x}_{k|k}^{t+1}-\nu^{t+1}$ (9c)
for $t=0,\dots,n_{a}-1$, where $\rho>0$ is a hyper-parameter to be calibrated and “prox” is the proximal operator [bib15]. As shown in [bib5], in the convex case, the ADMM iterations (9a)–(9c) converge to the optimizer of (8) as $n_{a}\rightarrow\infty$, and often converge to a solution of acceptable accuracy within a few tens of iterations. Iteration (9c) is straightforward to compute; iteration (9b) can be solved explicitly and efficiently with complexity $\mathcal{O}(n_{x})$ for a wide range of non-smooth and non-convex regularization functions $g$, such as $g(x)=\|x\|_{0}$, $g(x)=\|x\|_{1}$, and the indicator function $g(x)=0$ if $x_{\rm min}\leq x\leq x_{\rm max}$ or $\infty$ otherwise [bib15]. Iteration (9a) can be rewritten as
$\hat{x}_{k|k}^{t+1}=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+\left\lVert\overline{y}_{k}^{m}-\overline{C}_{k}x_{k}\right\rVert^{2}_{\overline{R}_{k}^{-1}}$ (10)
where $\overline{y}_{k}^{m}=[(y_{k}^{m})^{\prime}\ (\nu^{t}-w^{t})^{\prime}]^{\prime}$, $\overline{C}_{k}=[C_{k}^{\prime}\ I]^{\prime}$, and $\overline{R}_{k}=\left[\begin{smallmatrix}R_{k}&0\\ 0&\rho^{-1}I\end{smallmatrix}\right]$. Therefore, iteration (9a) can be performed directly in the correction step of the EKF by including $n_{x}$ additional “fake” state measurements $\nu^{t}-w^{t}$ with covariance matrix $\rho^{-1}I$.
Algorithm 1 summarizes the proposed extension of EKF with ADMM iterations (EKF-ADMM). The algorithm returns the estimate $\hat{x}_{k|k}$ of the parameter vector $x$ obtained after processing NN measurements. It also returns the last value of $\nu$, which could be used as an alternative estimate of $x$ too; for example, in case $g$ is the indicator function of a constraint set, $\nu$ would be guaranteed to be feasible. Note that the dual vector $w$ is not reset at each EKF iteration $k$; it is used as a warm start for the next $n_{a}$ ADMM iterations at step $k+1$, as the solutions $\hat{x}_{k|k}$ at consecutive time instants $k$ are usually similar.
3.1 Computational complexity
Given the block-diagonal structure of the measurement noise covariance matrix $\overline{R}_{k}$, (10) can be rewritten as
$\hat{x}_{k|k}^{t+1}$ $=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+$ (11)
$+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+\rho\sum_{i=1}^{n_{x}}\left\lVert x_{k,i}-\nu^{t}_{i}+w^{t}_{i}\right\rVert_{2}^{2}$
which highlights the separation of the contributions of the true measurements $y_{k}^{m}$ and of the fake regularization measurements $\nu^{t}-w^{t}$ in the correction step; moreover, we can process the measurements $\nu^{t}_{i}-w^{t}_{i}$ separately one by one. This allows designing a computationally more efficient and numerically robust version of the proposed EKF-ADMM algorithm, as the correction due to the true measurements $y_{k}^{m}$ can be performed only once, instead of $n_{a}$ times, as it does not change with $t$. Moreover, there is no need for any matrix inversion when processing the fake measurements, as by processing them one by one, the matrix inversion required to compute the Kalman gain becomes a simple division, since each measurement is just a scalar value. Assuming a complexity $\mathcal{O}(n_{x})$ for evaluating the proximal operator, EKF-ADMM has complexity $\mathcal{O}(n_{x}^{3}+n_{a}n_{x}^{2})$, which is the same order of the full EKF for general state estimation. Moreover, EKF-ADMM has the same number of Jacobian matrices evaluations than the classical EKF, which is usually the most time-consuming part in case $x$ represents the weights and bias terms of a neural network model to learn. Summarizing, the proposed approach is computationally efficient and, if the Kalman filter is implemented using numerically robust factored or square-root modifications [bib16], the method is appealing for embedded applications.
Algorithm 1 EKF-ADMM
$0\hat{x}_{0|-1},P_{0|-1}^{-1},\nu=\hat{x}_{0|-1},w=0,\rho>$0
for $k=0,\dots,N-1$ do
  $K_{k}=P_{k|k-1}\overline{C}_{k}^{T}(\overline{R}_{k}+\overline{C}_{k}P_{k|k-1}\overline{C}_{k}^{T})^{-1}$
  for $t=0,\dots,n_{a}-1$ do
   $\hat{x}_{k|k}\leftarrow\hat{x}_{k|k-1}+K_{k}\bigg(\begin{pmatrix}y_{k}^{m}\\ \nu-w\end{pmatrix}-\overline{C}_{k}\hat{x}_{k|k-1}\bigg)$
   $\nu\leftarrow\text{prox}_{\frac{g}{\rho}}(\hat{x}_{k|k}+w)$
   $w\leftarrow w+\hat{x}_{k|k}-\nu$
  end for
  $P_{k|k}=(I-K_{k}\overline{C}_{k})P_{k|k-1}$
  $\hat{x}_{k+1|k}=\hat{x}_{k|k}$
  $P_{k+1|k}=P_{k|k}+Q_{k}$
end for
return $\hat{x}_{N-1|N-1}$, $\nu$
4 Regret analysis
We investigate the theoretical properties of EKF-ADMM for linear time-varying models, i.e., models of the form
$y_{k}=h_{k}(x_{k})=C_{k}x_{k}$ (12)
where $C_{k}$ are now given time-varying matrices for $k=0,1,\ldots,N-1$, and convex regularization terms $g$. In particular, we want to evaluate the ability of the algorithm to solve the optimization problem $\min_{x}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$ online, where $f_{k}(x)=\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x\right\rVert^{2}_{R_{k}^{-1}}$, via the following two regret functions $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(\nu_{k}))-\min_{x,\nu\in\mathcal{S}}\sum_{k=0}^{N-1}(f_{k}(x)+g(\nu))$ and $R_{c}(N)=\sum_{k=0}^{N-1}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}$, where, to simplify the notation, we have defined $x_{k}=\hat{x}_{k|k-1},P_{k}=P_{k|k-1},\forall k=0,1,\ldots,N-1$. Notice that $R_{f}(N)$ quantifies the loss we suffer by learning the model online instead of solving it in a batch way given all NN measurements, while $R_{c}(N)$ quantifies the violation of the constraint $x=\nu$. To ensure a proper behavior of EKF-ADMM, we want to prove a sublinear regret bound for both, i.e., $R_{f}(N)\leq\mathcal{O}(\sqrt{N})$ and $R_{c}(N)\leq\mathcal{O}(\sqrt{N})$ [bib2].
EKF-ADMM is a generalization of the online ADMM method proposed in [bib2], in which a sublinear regret bound is derived for the case $n_{a}=1$ and $P_{k}^{-1}=P^{-1}\succ 0,\forall k$, while, more recently, in [bib18] a sublinear regret bound has been derived for the case $n_{a}=1$ and $P_{k}^{-1}\succeq P_{k+1}^{-1},\forall k$. Here we will provide a sublinear regret in the case $n_{a}=1$ and $P_{k}^{-1}=P_{k+1}^{-1},\forall k\geq k_{n}\ll N$, which is a reasonable assumption as the EKF covariance matrix, when estimating the parameters of a model, usually has a transient and then reaches a steady-state value. By assuming $n_{a}=1$ and $P_{k}^{-1}=P_{k+1}^{-1},\forall k\geq k_{n}\ll N$, Algorithm 1 can be equivalently rewritten as in Algorithm 2.
Algorithm 2 EKF-ADMM ($n_{a}=1$, frozen $P$)
$0x_{0},P_{0}^{-1},\nu_{1}=x_{0},w_{0}=0,\rho,\eta>$0, $k_{n}\geq 0$
for $k=0,\dots,N-1$ do
  $\begin{aligned} x_{k+1}\leftarrow&\operatorname*{arg\,min}_{x}\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x\right\rVert^{2}_{R_{k}^{-1}}+w_{k}^{T}(x-\nu_{k})+\\ &+\frac{\rho}{2}\left\lVert x-\nu_{k}\right\rVert^{2}+\frac{\eta}{2}\left\lVert x-x_{k}\right\rVert^{2}_{P_{k}^{-1}}\end{aligned}$
  $\begin{aligned} \nu_{k+1}\leftarrow&\operatorname*{arg\,min}_{\nu}g(\nu)+w_{k}^{T}(x_{k+1}-\nu)+\\ &+\frac{\rho}{2}\left\lVert x_{k+1}-\nu\right\rVert^{2}=\text{prox}_{\frac{g}{\rho}}(x_{k+1}+\frac{w_{k}}{\rho})\end{aligned}$
  $w_{k+1}\leftarrow w_{k}+\rho(x_{k+1}-\nu_{k+1})$
  if $k<k_{n}$ then
   $P_{k+1}^{-1}\leftarrow(Q_{k}+(P_{k}^{-1}+\overline{C}_{k}^{T}\overline{R}_{k}^{-1}\overline{C}_{k})^{-1})^{-1}$
  else
   $P_{k+1}^{-1}\leftarrow P_{k}^{-1}$
  end
end for
return $x_{N},\nu_{N}$
The following Theorem 4.1 is an extension of [bib2, Theorem 4], and provides conditions for sublinear regret bounds of Algorithm 2 in the case of a linear time-varying model (12) and convex regularization function $g$.
Theorem 4.1.
Let $\{x_{k},\nu_{k},w_{k}\}_{k=0}^{N-1}$ be the sequence generated by Algorithm 2 and let $x^{\star},\nu^{\star}$ be the best solution in hindsight, i.e. $x^{\star},\nu^{\star}=\operatorname*{arg\,min}_{x,\nu\in\mathcal{S}}\sum_{k=0}^{N-1}(f_{k}(x)+g(\nu))$. Let the following assumptions hold:
  1. A1.
    $\exists\alpha,G_{f},D_{x},D_{\nu},F>0$ such that $\forall k=0,\dots,N-1$:
    1. (a)
      $\left\lVert x-y\right\rVert^{2}_{P_{k}^{-1}}\geq\alpha\left\lVert x-y\right\rVert_{2}^{2}$, $\forall x,y$
    2. (b)
      $\left\lVert\nabla f_{k}(x_{k})\right\rVert_{2}^{2}=\left\lVert C_{k}^{T}R_{k}^{-1}(C_{k}x_{k}-y_{k}^{m})\right\rVert_{2}^{2}\leq G_{f}^{2}$
    3. (c)
      $\frac{1}{2}\left\lVert x^{\star}\right\rVert^{2}_{P_{k}^{-1}}\leq D_{x}^{2}$ and $\left\lVert\nu^{\star}\right\rVert^{2}_{2}\leq D_{\nu}$
    4. (d)
      $f_{k}(x_{k+1})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star}))\geq-F$
  2. A2.
    $\exists M_{k_{n}}\geq 0$ such that $\frac{1}{2}\sum_{k=1}^{k_{n}}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}\leq M_{k_{n}}$
  3. A3.
    To ease the notation, $x_{0}=0,g(0)=0$ and $g(\nu)\geq 0$.
Then, if $\eta=\frac{G_{f}\sqrt{N}}{D_{x}\sqrt{2\alpha}}$ and $\rho=\sqrt{N}$, the following sublinear regret bounds are guaranteed:
$R_{f}(N)\leq\frac{\sqrt{N}D_{\nu}}{2}+\frac{G_{f}D_{x}\sqrt{N}}{\sqrt{2\alpha}}+\frac{G_{f}\sqrt{N}(D_{x}^{2}+M_{k_{n}})}{D_{x}\sqrt{2\alpha}}$ (13a)
$R_{c}(N)\leq 2F\sqrt{N}+D_{\nu}+\frac{2G_{f}}{D_{x}\sqrt{2\alpha}}(D_{x}^{2}+M_{k_{n}}).$ (13b)
Proof. See Appendix A. $\Box$
Corollary 4.2.
Consider the linear time-invariant case $C_{k}\equiv C_{0}$, $\forall k\geq 0$. If the steady-state Kalman filter is used, then Theorem 4.1 holds with $M_{k_{n}}=0$.
In general, as proved in [bib18], Theorem 4.1 holds with $M_{k_{n}}=0$ whenever $P_{k}^{-1}\succeq P_{k+1}^{-1},\forall k$. Intuitively, this means that for online model adaptation we need to limit the importance of the previous samples to promptly adapt the model to changes and therefore bound the regret function. This can be accomplished, for example, using the EKF with a proper forgetting factor [bib3].
5 Simulation results
We evaluate the performance of the proposed EKF-ADMM algorithm on three different examples: online LASSO [bib22], online training of a neural network on data from a static model under $\ell_{1}$ regularization or bound constraints, online adaptation of a neural network on data from a time-varying model under $\ell_{0}$ regularization.
5.1 Online LASSO
Consider the LASSO problem $\min_{x}\sum_{k=0}^{N-1}(\frac{1}{2}\left\|y_{k}^{m}\right$. - $\left.C_{k}x\right\|^{2}_{R^{-1}}$ ++ $\lambda\left\lVert x\right\rVert_{1})$, where $x\in\mathbb{R}^{3}$ is the parameter vector, $C_{k}\in\mathbb{R}^{2\times 3}$ are randomly generated matrices with $C_{k,ij}\sim\mathcal{N}(0,1)$, $y_{k}^{m}=C_{k}x_{\rm true}+r_{k}\in\mathbb{R}^{2}$ is the vector of measurements and $r_{k}$ is random measurement noise, $r_{k,i}\sim{\mathcal{N}}(0,10^{-2})$. We will evaluate the behavior of the regret functions $R_{f}(N)$ and $R_{c}(N)$ as $N\rightarrow\infty$ when using Algorithm 2. The following settings are used: $P_{0|-1}=I,Q_{k}=10^{-6}I,R=10^{-3}I,k_{n}=10^{3},\rho=10^{4}\sqrt{N}$ and $\eta=10^{-6}\sqrt{N}$, where $P_{0|-1},Q_{k}$ and $R$ are manually tuned to maximize the Kalman filter performance without regularization and $\rho,\eta$ are chosen according to the expressions provided in Theorem 4.1. Results for different values of $\lambda$ are shown in Figure 1. In this case, Theorem 2 holds and, as expected, both the regrets $R_{f}(N)$ and $R_{c}(N)$ decrease as the number NN of samples increases.
Figure 1: Objective and constraint regret for online LASSO.
5.2 Online learning of a static model
Consider the dataset generated by the static nonlinear model $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{10}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$. We want to train online a neural network $h_{k}(x)$ with 22 layers, 88 neurons in each layer, and $\tanh$ activation function, with $n_{x}=105$ trainable weights in total. The training is performed on $N=10^{5}$ randomly generated data points, where $z_{k,i}\sim\mathcal{U}_{[-10,10]}$ and $r_{k}\sim{\mathcal{N}}(0,10^{-2})$. Let $\{x_{k}\}_{k=0}^{N-1}$ be the sequence of weights generated by Algorithm 1: we evaluate the online adaptation performance by means of the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{x}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$, where $f_{k}(x)=\frac{1}{2}\left\lVert y_{k}^{m}-h_{k}(x)\right\rVert^{2}$, and the quality of a given solution $x$ using the performance indices $\text{Loss}(x)=\frac{1}{N}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$, $\text{Mse}(x)=\frac{1}{N}\sum_{k=0}^{N-1}f_{k}$, $\text{Reg}(x)=\frac{1}{N}\sum_{k=0}^{N-1}g(x)$ and $\text{Cv}(x)=\left\lVert x-\Pi_{\mathcal{C}}(x)\right\rVert_{2}^{2}$, where, given a constraint set $\mathcal{C}\subseteq\mathbb{R}^{n_{x}}$, $\Pi_{\mathcal{C}}(x)$ is the projection of the point $x$ onto $\mathcal{C}$. The training is performed in MATLAB R2022a on an Intel Core i7 12700H CPU with 16 GB of RAM, using the library CasADi [AGHRD19] to compute the required Jacobian matrices via automatic differentiation. All results are averaged over 20 runs starting from different initial conditions, that were randomly generated using the well known Xavier’s initialization strategy.
5.2.1 $\ell_{1}$ regularization
We train the neural model under the regularization function $g(x)=\lambda\left\lVert x\right\rVert_{1}$, with $\lambda=10^{-4}$. We selected the following hyper-parameters: $\rho=10\lambda$, $n_{a}=1$, $Q_{k}=10^{-4}I$, $R_{k}=I$ and $P_{0|0}=100I$. We compare the results to different online optimization alternatives: online ADMM [bib2] with constant matrix $P=10^{-2}I$ (online-ADMM), EKF-ADMM with time-varying $\rho_{k}=10^{\frac{k}{N}-2}\lambda$ (EKF-ADMMtv), EKF with $\ell_{1}$-regularization [bib1] (EKF-$\ell_{1}$) and SMIDAS [bib24] with learning rate $\eta=5\cdot 10^{-2}$. Notice that $P_{0|0},R_{k}$ and $Q_{k}$ are chosen by manually tuning EKF-$\ell_{1}$ and then used for all other approaches, while all the remaining method-specific hyperparameters have been manually tuned to maximize their performance. The reason for choosing a time-varying $\rho_{k}$ is that fake measurements are usually not accurate initially, so that it is better to start with a higher value of $\frac{1}{\rho_{k}}$ and then decrease it progressively. In addition, we compare with two offline batch algorithms: NAILM [bib4] and LBFGS [bib23], the latter using the Python library jax-sysid. Note that all the online algorithms consume the dataset only once (1 epoch), except NAILM and LBFGS that run over 150 and 5000 epochs respectively. The results obtained at the end of the training phase are reported in Table 1.
Table 1: Online learning a static model of (1) with $\ell_{1}$ regularization: mean (standard deviation) Loss, Mse, sparsity ratio and execution time obtained over 20 runs.
Loss ($10^{-3}$) Mse ($10^{-3}$) Sparsity (%) Time [s]
LBFGS [bib23] $5.40$ ($0.72$) $1.03$ ($0.19$) $80.66$ ($5.32$) $80.51$ ($2.42$)
NAILM [bib4] $5.24$ ($0.48$) $1.06$ ($0.15$) $63.85$ ($5.00$) $235.41$ ($52.78$)
EKF-ADMM $5.99$ ($0.68$) $1.44$ ($0.17$) $45.28$ ($4.98$) $58.27$ ($1.41$)
EKF-ADMMtv $5.27$ ($0.46$) $1.29$ ($0.42$) $57.00$ ($7.95$) $55.90$ ($1.41$)
online-ADMM [bib2] $10.38$ ($1.7$) $4.68$ ($1.8$) $3.62$ ($2.21$) $530.69$ ($29.09$)
EKF-$\ell_{1}$ [bib1] $5.47$ ($0.67$) $1.42$ ($0.26$) $56.42$ ($7.67$) $12.46$ ($0.27$)
SMIDAS [bib24] $89.93$ ($36.3$) $84.67$ ($36.6$) $61.71$ ($9.98$) $4.11$ ($1.11$)
Among the online approaches, EKF-ADMMtv provides the lowest loss: this is also true during the training phase, as shown in Figure 2. The online learning performance of EKF-ADMM can be also evaluated by looking at the regret function in Figure 3, where it is also apparent that, compared to the other approaches, the proposed algorithm improves the solution quality faster.
Figure 2: Online learning with $\ell_{1}$ regularization: Loss, Mse and Reg averaged over 2020 runs.
Figure 3: Online learning with $\ell_{1}$ regularization: regret and sample regret averaged over 2020 runs.
5.2.2 Bound constraints
Let us now repeat the training under the bound constraints imposed by the regularization function $g(x)=0$ if $x\in\mathcal{C}$ and $g(x)=+\infty$ otherwise, where $\mathcal{C}=\{x\in\mathbb{R}^{n_{x}}:\ |x_{i}|\leq 0.5\}$. We use the hyper-parameters $\rho=1,n_{a}=5,Q_{k}=10^{-4}I,P_{0|0}=100I$, and $R_{k}=I$. In this example, we also compare with a simple clipping step of the Kalman filter (EKF-CLIP) and an online projected gradient method (OGD-proj) [bib7] with learning rate $\eta=10^{-3}$. Matrices $P_{0|0},R_{k}$ and $Q_{k}$ are the same as for the $\ell_{1}$ example and are in common for all the Kalman-like approaches, while all the other hyperparameters have been manually re-tuned to optimize performance in the bounds constraints example. Results obtained at the end of the training phase are reported in Table 2. Among the online approaches, considering the final Mse, Cv and execution time, EKF-ADMM provides the best quality solution, while OGD-proj attains similar performance but at a slower pace. Figure 4 shows the performance of the solution during the training phase.
Table 2: Online learning a static model (1) with bound constraints: mean (standard deviation) Mse, constraints violation, and execution time obtained over 20 runs.
Mse Cv ($10^{-6}$) Time [s]
LBFGS [bib23] $0.122$ ($0.011$) 0 (0) $75.87$ ($5.58$)
NAILM [bib4] $0.137$ ($0.013$) $0.38$ ($0.71$) $101.82$ ($3.88$)
EKF-ADMM $0.131$ ($0.011$) $10.76$ ($4.93$) $70.46$ ($3.45$)
online-ADMM [bib2] $0.129$ ($0.010$) $90.77$ ($59.37$) $610.73$ ($9.82$)
EKF-CLIP $0.214$ ($0.048$) 0 (0) $11.89$ ($0.12$)
OGD-proj [bib7] $0.137$ ($0.013$) 0 (0) $2.56$ ($0.15$)
Figure 4: Online learning with bounds: Mse and constraints violation averaged over 20 runs.
5.3 Online learning of a time-varying model
We test now the ability of EKF-ADMM to adapt the same neural network model, under $\ell_{0}$ regularization, when the data-generating system switches as follows: $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{10}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $k\leq\frac{N}{3}$, $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{2}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $\frac{N}{3}<k\leq\frac{2N}{3}$ and $y^{m}_{k}=\frac{0.3\cdot z_{k,1}^{2}-e^{\frac{z_{k,2}}{2}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $\frac{2N}{3}<k$, with $N=1.5\cdot 10^{5}$, $z_{k,i}\sim\mathcal{U}_{[-10,10]}$ and $r_{k}\sim{\mathcal{N}}(0,10^{-2})$. We evaluate the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{z_{1},z_{2},z_{3}}\sum_{i=1}^{3}r_{i}(z_{i})$, with $r_{i}(z_{i})=\sum_{k=(i-1)\frac{N}{3}}^{i\frac{N}{3}}(f_{k}(z_{i})+g(z_{i}))$, where $\{x_{k}\}_{k=0}^{N-1}$ is the sequence generated by Algorithm 1. The regularization term is $g(x)=\lambda\left\lVert w\right\rVert_{0}$, with $\lambda=10^{-4}$, and we use the EKF-ADMM hyper-parameters $\rho=10^{3}\cdot\lambda$, $p_{a}=1$, $Q_{k}=10^{-4}I$, and $P_{0|0}=100I$, which were manually tuned by optimizing the EKF performance. Since the model is now time-varying, we will also use an EKF implementation with forgetting factor $\alpha=0.9$ [bib3]. The resulting regret function is shown in Figure 5. It is apparent that EKF-ADMM can effectively track changes of the underlying data-generating system.
We have noticed that the choice of $\rho$ is crucial to obtain good adaptation performance. While for some values of $\rho$ the convergence is slow and not satisfactory, with limited effort, in all the reported examples, we could find appropriate values that gave good performance. We remark that, in the general nonlinear case, this might not always be the case, as both the EKF and ADMM are not guaranteed to converge, which might complicate tuning $\rho$. This could limit the use of ADMM iterations within the EKF framework in certain challenging applications.
Figure 5: Online learning with $\ell_{0}$ regularization: regret and sample regret averaged over 2020 runs.
6 Conclusions
We have proposed a novel algorithm for online learning of nonlinear parametric models under non-smooth regularization using a combination of EKF and ADMM, for which we derived a sublinear regret bound for the convex linear time-varying case. The approach is computationally cheap and is suitable for factorized or square-root implementations that can make it numerically robust, and is therefore very appealing for embedded applications of adaptive control, such as adaptive model predictive control. The effectiveness of the approach has been evaluated in three numerical examples. Future investigations will focus on extending the approach to the recursive identification of parametric nonlinear state-space dynamical models from input/output data under non-smooth regularization, in which both the hidden states and the parameters are jointly estimated.
Appendix A Proof of Theorem 1
Starting from $R_{f}(N)$, since $x_{k+1},\nu_{k+1}$ are the optimal solutions of the first two optimization problems in Algorithm 2 and since $w_{k}=w_{k+1}-\rho(x_{k+1}-\nu_{k+1})$ we have that $\nabla f_{k}(x_{k+1})=-(w_{k+1}-\rho(\nu_{k}-\nu_{k+1}))-\eta(P_{k}^{-1}x_{k+1}-P_{k}^{-1}x_{k})$ and $w_{k+1}\in\partial g(\nu_{k+1})$, where $\partial g$ is the subgradient of $g$. Due to the convexity of $f_{k}(\cdot)$ and $g(\cdot)$, $f_{k}(x_{k+1})-f_{k}(x^{\star})\leq\nabla f_{k}(x_{k+1})^{T}(x_{k+1}-x^{\star})=-w_{k+1}^{T}(x_{k+1}-\nu^{\star})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2}+\left\lVert x_{k+1}-\nu_{k+1}\right\rVert^{2}-\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2})+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}$- $\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}})$, where the equality derives from a succession of square completions and by recalling that $x^{\star}=\nu^{\star}$, and $g(\nu_{k+1})-g(\nu^{\star})\leq w_{k+1}^{T}(\nu_{k+1}-\nu^{\star})$. Summing the two inequalities together and noticing that $w_{k+1}^{T}(x_{k+1}-\nu_{k+1})+\frac{\rho}{2}\lVert x_{k+1}-\nu_{k+1}\rVert^{2}=\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})$, we obtain:
$f_{k}(x_{k+1})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star}))\leq$ (14)
$\leq\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})-\frac{\rho}{2}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}+$
$\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}$
$-\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}})$
Considering that $f_{k}(x_{k})-f_{k}(x_{k+1})\leq\nabla f_{k}(x_{k})^{T}(x_{k}-x_{k+1})\leq\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\alpha\eta}{2}\left\lVert x_{k}-x_{k+1}\right\rVert^{2}$, where the second inequality is due to Fenchel-Young’s inequality, and considering Assumption A1.a of the theorem, we have that $f_{k}(x_{k})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star}))\leq\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})-\frac{\rho}{2}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2})$. Summing from 0 to N1N-1 and considering Assumption A3 we get
$R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star})))$
$+g(\nu_{0})-g(\nu_{N})\leq\frac{1}{2\rho}(\left\lVert w_{0}\right\rVert^{2}-\left\lVert w_{N}\right\rVert^{2})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{0}\right\rVert^{2}$
$-\left\lVert\nu^{\star}-\nu_{N}\right\rVert^{2})+\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}-x_{0}\right\rVert^{2}_{P_{0}^{-1}}$
$+\frac{\eta}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}$
and, therefore, $R_{f}(N)\leq\frac{\rho}{2}\left\lVert\nu^{\star}\right\rVert^{2}$ $\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}\right\rVert^{2}_{P_{0}^{-1}}+\frac{\eta}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}$. Because of Assumption A2, $\frac{1}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}=\frac{1}{2}\sum_{k=1}^{k_{n}}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}\leq M_{k_{n}}$, and taking into account Assumptions A1.b and A1.c we get $R_{f}(N)\leq\frac{\rho D_{\nu}}{2}+\frac{NG_{f}^{2}}{2\alpha\eta}+\eta(D_{x}^{2}+M_{k_{n}})$. Setting $\eta\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{G_{f}\sqrt{N}}{D_{x}\sqrt{2\alpha}}$ and $\rho\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{N}$, we get the sublinear regret bound $R_{f}(N)\leq\frac{\sqrt{N}D_{\nu}}{2}+\frac{G_{f}D_{x}\sqrt{N}}{\sqrt{2\alpha}}+\frac{G_{f}\sqrt{N}(D_{x}^{2}+M_{k_{n}})}{D_{x}\sqrt{2\alpha}}$. Considering now $R_{c}(N)$, we can rearrange (14) and consider Assumption A1.d, $\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}\leq\frac{2F}{\rho}+\frac{1}{\rho^{2}}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})+(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})++\frac{\eta}{\rho}\left(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x_{k+1}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}\right).$ Summing from 0 to N1N-1, we get:
$R_{c}(N)=\sum_{k=0}^{N-1}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}\leq\frac{2FN}{\rho}+\left\lVert\nu^{\star}\right\rVert^{2}$
[…]
Considering Assumptions A1.c and A2, we have $R_{c}(N)\leq\frac{2FN}{\rho}+D_{\nu}+\frac{2\eta}{\rho}(D_{x}^{2}+M_{k_{n}})$ and setting $\eta\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{G_{f}\sqrt{N}}{D_{x}\sqrt{2\alpha}}$ and $\rho\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{N}$ we finally get $R_{c}(N)\leq 2F\sqrt{N}+D_{\nu}+\frac{2G_{f}}{D_{x}\sqrt{2\alpha}}(D_{x}^{2}+M_{k_{n}})$. $\Box$